# GeoVLMath: Enhancing Geometry Reasoning in Vision-Language Models via Cross-Modal Reward for Auxiliary Line Creation

### v1

@@ Abstract
Auxiliary lines are essential for solving complex geometric problems but remain challenging for large vision-language models (LVLMs). Rather than editing diagrams to draw auxiliary lines, which current image editing models struggle to render with geometric precision, we generate textual descriptions of auxiliary-line constructions to better align with the representational strengths of LVLMs. To bridge the gap between textual descriptions and spatial structure, we propose a reinforcement learning framework that enhances diagram-text alignment. At the core of our approach is a cross-modal reward that evaluates how well the generated auxiliary-line description for an original diagram matches a ground-truth auxiliary-line diagram. Built on this reward, we present GeoVLMath, an open-source LVLM tailored to auxiliary-line reasoning in solid geometry. This fine-grained signal drives a GRPO-based RL stage, yielding precise diagram-text alignment. To support training, we develop a scalable data creation pipeline and construct AuxSolidMath [^1] The dataset is available at https://huggingface.co/datasets/shasha/AuxSolidMath, a dataset of 3,018 real-exam geometry problems with paired diagrams and aligned textual fields. At the 3B and 7B scales, GeoVLMath achieves competitive and often superior performance compared with strong open-source and proprietary LVLMs on auxiliary-line reasoning benchmarks.
@@ 1 Introduction
Geometric problems constitute an important category of mathematical tasks, characterized by intricate spatial structures and multi-step reasoning processes (Ma et al., 2024). They are commonly divided into plane geometry and solid geometry. This study focuses on solid geometry, where reasoning over three-dimensional spatial relations is substantially more complex. Such problems rarely yield to direct application of standard theorems; instead, they often require the deliberate introduction of auxiliary lines [^2] In this paper, we use auxiliary lines broadly to include both additional lines and coordinate systems. to reveal hidden geometric structure and enable further analysis. These constructions are essential for anchoring visual diagrams to formal symbolic reasoning and for providing the intermediate steps required for rigorous problem solving.
To validate the above idea, we conduct a pilot study with three preliminary experiments: one with incorrect auxiliary lines (Incorrect Aux), one without auxiliary lines (No Aux), and one with correct auxiliary lines (Correct Aux). As shown in Figure 1, the use of correct auxiliary lines achieves the highest accuracy, whereas incorrect auxiliary lines lead to the poorest performance. One possible explanation is that inaccurate auxiliary lines tend to misdirect reasoning and produce errors, while precise auxiliary lines uncover key spatial relationships, thereby enhancing solution accuracy.
Given that accurate auxiliary lines are crucial for solution correctness, the key question is how to obtain them reliably. The most straightforward approach is to draw auxiliary lines directly on the original diagram. However, current image editing models cannot accurately draw auxiliary lines (as illustrated in Figure 5). Another representative approach adopts a tool-use pipeline, exemplified by Visual Sketchpad (Hu et al., 2024): an LVLM is prompted to generate code that draws auxiliary lines, which is then rendered into an augmented diagram. Although effective in controlled settings, this method faces two key limitations. First, it depends on the specific coordinate position of the diagram elements, which is rarely available in real-world problems.
@@
Figure 1: Pass@1 of Qwen2.5-VL-72B-Instruct and Gemini-2.5-Flash. “Aux” denotes auxiliary-line description.
Additionally, it relies on the LVLM’s ability to generate highly accurate code for auxiliary lines. This dependency constrains its robustness and limits its applicability to broader classes of geometric problems. Taken together, these limitations indicate that neither direct image editing nor coordinate-dependent tool usage approaches can reliably produce accurate auxiliary lines in current technology, motivating an approach that avoids image manipulation and strict coordinate requirements.
@@
Motivated by the success of textual chain-of-thought, we avoid image editing and explicit coordinates by expressing auxiliary-line constructions in text and verifying them against the diagram. Our main idea is to design a cross-modal reward model that measures the consistency between a generated textual auxiliary-line description for the original diagram and a ground-truth auxiliary-line diagram. The reward is computed by jointly encoding the original diagram and the generated auxiliary-line description, and then comparing this pair with the ground-truth auxiliary-line diagram, providing geometry-aware supervision without requiring coordinate assumptions or image manipulation. Building on this reward signal, we train a policy model using Group Relative Policy Optimization (GRPO) to obtain geometry-consistent, generalizable constructions. Training follows a two-stage paradigm inspired by recent progress in reinforcement learning (RL) for reasoning (e.g., DeepSeek-R1 (Guo et al., 2025)): supervised fine tuning (SFT) for cold start, followed by GRPO-based RL to further elicit structured reasoning and strengthen diagram-text alignment. We instantiate the framework as GeoVLMath (3B/7B), a vision-language model tailored to auxiliary-line-based geometric reasoning. Through the cross-modal supervision, GeoVLMath achieves strong alignment between text and geometric structure, enabling faithful reasoning on complex diagrams.
@@
To effectively train the above model, we require a high-quality dataset that captures both visual and symbolic aspects of real-world geometry problems. Yet, creating such a dataset is inherently challenging due to the need for automation, scalability, and semantic precision across diverse and noisy educational materials. In response to these challenges, we develop a robust and scalable data construction pipeline specifically designed to transform raw high school exam papers into structured multimodal instances suitable for training LVLMs. (1) First, we develop automated scripts to identify and filter geometry problems that satisfy predefined criteria from a large corpus of exam papers; (2) Next, we apply a fully automated pipeline to deduplicate problems and extract their associated geometric diagrams, which are filtered to remove low-resolution or unclear diagrams and then saved for subsequent processing; (3) Subsequently, we train a specialized text extraction model capable of handling MathType equations. This model accurately parses the problem description, the final answer, and the auxiliary-line description from exam papers and converts them into a structured JSON format; (4) Finally, we perform a manual verification step to ensure the accuracy, completeness, uniqueness, and semantic consistency of each data instance. While the pipeline is largely automated, this lightweight manual verification step is essential for maintaining data quality, particularly when handling complex symbolic expressions and diagrammatic content in real-world settings. Based on this pipeline, we construct AuxSolidMath, a curated dataset of 3,018 solid geometry problems in a rich multimodal format, comprising the problem description, the final answer, the auxiliary-line description, the original diagram, and the auxiliary-line diagram. To our knowledge, AuxSolidMath is the first systematically constructed dataset explicitly tailored to auxiliary-line-based solid geometry reasoning. Beyond this specific resource, our pipeline offers a scalable framework that can be readily adapted to generate large-scale, high-quality datasets for multimodal reasoning tasks.
@@
To assess the effectiveness of our model, GeoVLMath, we conduct a comprehensive evaluation against frontier open-source and closed-source LVLMs. Despite its relatively modest parameter scale, GeoVLMath achieves highly competitive performance, frequently outperforming substantially larger models such as Qwen2.5-VL-32B-Instruct (Bai et al., 2025) and GPT-4o (Hurst et al., 2024). These results suggest that supervision grounded in auxiliary-line constructions is more effective for enhancing geometric reasoning than simply scaling model parameters. Furthermore, our evaluation protocol highlights the utility of datasets augmented with auxiliary lines in revealing the limitations of existing LVLMs in structured visual reasoning, and provides a principled foundation for future model design and evaluation.
The main contributions of this work are threefold:
  • Cross-modal reward for auxiliary-line alignment. We propose a geometry-aware scalar reward that contrasts a fused diagram-text representation with the ground-truth auxiliary-line diagram, requiring no coordinates or image edits. Training with GRPO turns this signal into effective RL supervision, improving diagram-text alignment and spatial reasoning.
  • AuxSolidMath: a dataset for auxiliary-line geometric reasoning. We design a scalable, modular pipeline to construct AuxSolidMath, a dataset tailored to auxiliary-line reasoning in solid geometry problems. Curated from authentic high-school exams, it comprises 3,018 richly annotated multimodal instances pairing diagrams with aligned textual fields, providing training data for our proposed framework. AuxSolidMath will be publicly released on Hugging Face to enable community use and further research.
  • GeoVLMath(3B/7B): Competitive Small-Scale Models for Geometric Reasoning. We introduce GeoVLMath, a LVLM trained on AuxSolidMath and optimized with a GRPO-based RL framework. Despite its relatively modest parameter scale, GeoVLMath achieves performance comparable to, or even surpassing, that of substantially larger models such as Qwen2.5-VL-32B-Instruct and GPT-4o on geometric reasoning tasks requiring auxiliary line constructions.
@@ 2.1 Problem Definition
We study geometry problems whose solution requires the active construction of auxiliary lines. Formally, each instance is represented as a pair $\langle I,q\rangle$, where $I$ denotes an original diagram, and $q$ is a question grounded in $I$. The goal is to generate a solution $y$ that includes the auxiliary lines auxaux, a sequence of deductive steps, and a final answer ansans. Since auxiliary lines are absent from the original diagram $I$, they must be constructed during reasoning. The deliberate introduction of these lines uncovers latent spatial relations and formalizes them as explicit geometric constraints, thereby enabling solutions that would otherwise be difficult to derive from the original diagram alone.
@@ 2.2 Framework Overview
We introduce a two-stage training framework for LVLMs that integrates the auxiliary-line construction into the reasoning process (Hong et al., 2021). In the first stage, we apply supervised fine-tuning (SFT) on automatically synthesized chain-of-thought (CoT) data with explicit auxiliary-line steps, enabling the model to actively construct auxiliary lines, thereby establishing a good initialization. In the second stage, we further use reinforcement learning (Xia et al., 2015) to encourage the model to construct auxiliary lines that faithfully reflect the geometry of the diagram, boosting the precision of the solution. At the core is a cross-modal reward model that provides fine-grained feedback by scoring the agreement between the original diagram plus the generated auxiliary-line description and a reference diagram annotated with the correct auxiliary line. Overall, this framework combines direct supervision with structured visual feedback, resulting in more reliable auxiliary-line constructions and stronger geometric problem-solving performance.
Figure 2: Overview of the cross-modal reward-driven RL. We first fine-tune a cross-modal reward model on curated high-quality data to evaluate the correctness of auxiliary-line constructions. During the RL phase, the reward model’s consistency score is combined with a final-answer accuracy reward to produce a composite signal that updates the policy via GRPO.
2.3 Supervised Fine-Tuning for Cold-Start Initialization
We start with SFT on COT exemplars to provide a cold-start initialization for subsequent reinforcement learning. Each training instance is represented as a triplet $\langle I,q,y\rangle$, where $I$ denotes an original geometric diagram, $q$ is a natural language question, and $y=(y_{1},...,y_{T})$ is a step-by-step solution that includes the construction of necessary auxiliary lines auxaux, intermediate deductions, and the final answer ansans. To facilitate structured supervision and prepare for reward modeling in the RL stage, we mark auxiliary lines with special tokens [AUX] and [/AUX] within $y$. These markers expose auxiliary-line steps as explicit supervision signals. We fine-tune the model using a standard next-token prediction objective. The SFT loss is defined as:
$\mathcal{L}_{\text{SFT}}=-\mathbb{E}_{(I,q,y)\sim\mathcal{D}}\left[\sum_{t}^{T}\log P_{\theta}(y_{t}\mid I,q,y_{<t})\right]$ (1)
where $\mathcal{D}$ is the training set, $\theta$ denotes the model parameters, $y_{t}$ indicates the $t$-th output token, and $y_{<t}$ comprises all previously generated output tokens.
2.4 Cross-Modal Reward-Driven Reinforcement Learning
@@
As shown in Figure 1, accurate auxiliary-line constructions improve reasoning success. The key challenge is how to integrate these constructions into the reasoning process. The most intuitive approach is to draw auxiliary lines via image editing models or adopt coordinate-dependent pipelines, but limited editing reliability and the need for precise, accessible coordinates often render both impractical. Motivated by the recent advances of textual COT supervision (Xu et al., 2024; Zhang et al., 2025), we represent auxiliary-line constructions in natural language and propose a cross-modal reward model that scores diagram-text alignment between the original diagram together with the textual description and a reference diagram annotated with the ground-truth auxiliary lines. This resulting signal is geometry-aware yet agnostic to image editing and strict coordinate assumptions, enabling scalability across diverse diagram styles. We integrate this cross-modal reward, alongside a final-answer reward, into a GRPO-based RL stage to align intermediate constructions with the diagram while maintaining final-answer accuracy. An overview of the stage is illustrated in Figure 2.
2.4.1 cross-modal reward model for diagram-text alignment
@@
Given an original diagram $I$, a textual description $d$ of auxiliary lines (either the correct $d_{\text{aux}}$ or a perturbed $d^{-}_{\text{aux}}$), and a reference diagram $I^{+}$ annotated with the ground-truth auxiliary lines, the reward model compares the relations induced by applying $d$ to $I$ against the additional geometric structures present in $I^{+}$ but absent from $I$. Beyond surface similarity, it evaluates diagram-text relational consistency, such as parallelism, orthogonality, and angle bisection, and scores whether these relations are satisfied in the reference diagram. Accordingly, the reward exhibits three desirable properties: (1) cross-modal relational alignment, measuring diagram-text consistency at the level of geometric relations, rather than lexical similarity; (2) Coordinate-free and render-free, requiring neither pixel-accurate coordinates nor explicitly drawing the specified auxiliary lines on the diagram, since the alignment is computed from diagram-text correspondence; and (3) Fine-grained, providing intermediate scores for partially correct yet meaningful constructions, thereby enabling precise credit assignment across multi-step reasoning. In summary, the cross-modal reward precisely measures diagram-text spatial consistency, capturing whether and to what extent the generated auxiliary lines match the intended geometric structure, without explicitly rendering on the diagram or relying on coordinate-dependent methods. Next, we explain how to construct diagram-text supervision automatically and at scale, and train the cross-modal reward model accordingly.
@@
Constructing Diagram–Text Supervision. Each training example is represented as $\langle I,d,I^{+},\hat{y}\rangle$, where $\hat{y}=(r,s)$ contains a brief justification $r$ and a consistency score $s\in[0,1]$ indicating how well $d$ aligns with $I^{+}$ given $I$. We construct this supervision dataset via a fully automated pipeline (see Figure 2). Starting from high-quality triplets $\langle I,d_{\text{aux}},I^{+}\rangle$ sourced from the dataset described in Section 3, we introduce rule-based perturbations that simulate common errors, including partial deletion, intersection alteration, incorrect line connections, adding irrelevant lines, unrelated auxiliary lines. Based on these templates, we then leverage a large language model to generate diverse and linguistically varied negatives $d^{-}_{\text{aux}}$ that are lexically plausible yet geometrically inconsistent with the intended construction in the reference diagram. To evaluate the consistency between each description $d\in\{d_{\text{aux}},d^{-}_{\text{aux}}\}$ and its target construction $I^{+}$, we employ an LVLM-as-a-Judge strategy. Specifically, the LVLM is prompted to assess diagram-text alignment for the pair $\langle I,d\rangle$ against $I^{+}$ and to output both a natural language rationale $r$ and a scalar score $s\in[0,1]$. This automated evaluation provides interpretable justifications and continuous alignment signals, enabling scalable, fine-grained supervision that spans faithful descriptions through adversarial counterexamples.
@@
Training the Cross-Modal Reward Model. Given the input triplet $\langle I,d,I^{+}\rangle$, the model outputs a structured prediction $\hat{y}=(r,s)$, where $r$ is a rationale and $s\in[0,1]$ is a consistency score. We train the model by maximizing the conditional likelihood of the serialized output:
@@
Through this training, we obtain a reward model that provides precise, interpretable feedback on diagram-text alignment between auxiliary-line descriptions and the reference diagram. This model serves as a key component of our RL framework, guiding the policy toward auxiliary-line constructions that are geometrically consistent and diagram-grounded.

### v2

@@ Abstract
Auxiliary lines are essential for solving complex geometric problems but remain challenging for large vision-language models (LVLMs). Recent attempts construct auxiliary lines via code-driven rendering, a strategy that relies on accurate and executable code generation to produce visual renderings of the auxiliary lines for subsequent reasoning. However, in complex solid geometry settings, such a strong dependence on precise specifications substantially restricts the robustness of this strategy. Alternatively, we turn to a simpler and more stable solution, representing auxiliary-line constructions as structured textual descriptions. To bridge the gap between textual descriptions and spatial structure, we propose a reinforcement learning framework that enhances diagram-text alignment. The core is a cross-modal reward model that evaluates how well the generated auxiliary-line description matches the ground-truth auxiliary-line diagram. The reward signal drives a GRPO-based RL stage to yield informative auxiliary-line descriptions for the reasoning. To support the training and evaluation, we develop a scalable data pipeline and construct AuxSolidMath [^1] The dataset is now publicly available at https://huggingface.co/datasets/shasha/AuxSolidMath, a dataset of 3,018 real-exam geometry problems with paired diagrams and aligned textual fields. Based on this framework, we derive GeoVLMath, an LVLM for solving complex solid geometry.
GeoVLMath: Enhancing Geometry Reasoning in Vision-Language Models via Cross-Modal Reward for Auxiliary Line Creation
Shasha Guo1, Liang Pang1, Xi Wang2, Yanling Wang3, Huawei Shen1, Jing Zhang2thanks: Corresponding Author 1Institute of Computing Technology, Chinese Academy of Sciences, China 2Renmin University of China, China 3Zhipu AI, China {guoshasha, pangliang, shenhuawei}@ict.ac.cn {wangxi2022, zhang-jing}@ruc.edu.cn, yanlingwang777@gmail.com
@@ 1 Introduction
Geometric problems constitute an important category of mathematical tasks, characterized by intricate spatial structures and multi-step reasoning processes (Ma et al., 2024). They are commonly divided into plane geometry and solid geometry. This study focuses on solid geometry, where reasoning over three-dimensional spatial relations is substantially more complex. Such problems rarely yield to direct application of standard theorems; instead, they often require the deliberate introduction of auxiliary lines [^2] In this paper, we use auxiliary lines broadly to include both additional lines and coordinate systems. to reveal hidden geometric structure and enable further analysis. These auxiliary lines are essential for anchoring visual diagrams to formal symbolic reasoning and for providing the intermediate steps for rigorous problem-solving.
@@
Figure 1: Pass@1 results for Qwen2.5-VL-72B-Instruct and Gemini-2.5-Flash. “Aux” indicates the use of auxiliary-line descriptions.
To validate the above idea, we conduct a pilot study comprising three experimental settings: Incorrect Aux (with incorrect auxiliary lines), No Aux (without auxiliary lines), and Correct Aux (with correct auxiliary lines). As shown in Figure 1, the use of correct auxiliary lines achieves the highest accuracy, whereas incorrect auxiliary lines lead to the poorest performance. One possible explanation is that inaccurate auxiliary lines tend to misdirect reasoning and produce errors, while precise auxiliary lines uncover key spatial relationships, thereby enhancing solution accuracy.
Given that accurate auxiliary lines are crucial for the correctness of the solution, the key question is how to obtain them reliably. A seemingly straightforward solution is to explicitly construct them on the diagram. Indeed, several representative methods, such as Visual Sketchpad (Hu et al., 2024), V-Thinker (Qiao et al., 2025), and CodePlot-CoT (Duan et al., 2025), follow a unified paradigm in which the model generates Python-based drawing code to render intermediate diagrams that are then fed back into the reasoning process. Despite its conceptual simplicity, this code-driven visual construction paradigm exhibits inherent limitations in solid geometry. It critically relies on precise coordinate information and highly accurate code generation, and becomes particularly fragile when auxiliary lines involve cross-plane relations or skew structures, where minor rendering inaccuracies can distort spatial constraints and mislead reasoning. Our empirical results corroborate this limitation, showing that explicitly rendered auxiliary-line diagrams underperform text-based auxiliary-line formulations (See Table 2).
@@
In light of these observations, we adopt a simpler and more stable formulation by representing auxiliary-line constructions as structured textual descriptions. Our main idea is to design a cross-modal reward model that measures the consistency between a generated textual auxiliary-line description for the original diagram and a ground-truth auxiliary-line diagram. The reward is computed by jointly encoding the original diagram and the generated auxiliary-line description, and then comparing this pair with the ground-truth auxiliary-line diagram, providing geometry-aware supervision without requiring coordinate assumptions or image manipulation. Building on this reward signal, we train a policy model using Group Relative Policy Optimization (GRPO) to obtain geometry-consistent, generalizable constructions. Training follows a two-stage paradigm inspired by recent progress in reinforcement learning (RL) for reasoning (e.g., DeepSeek-R1 (Guo et al., 2025)): supervised fine-tuning (SFT) for cold start, followed by GRPO-based RL to further elicit structured reasoning and strengthen diagram-text alignment. We instantiate the framework as GeoVLMath, a vision-language model tailored to auxiliary-line-based geometric reasoning. Through the cross-modal supervision, GeoVLMath achieves strong alignment between text and geometric structure, enabling faithful reasoning on complex diagrams.
@@
To effectively train the above model, we require a high-quality dataset that captures both visual and symbolic aspects of real-world geometry problems. However, creating such a dataset is inherently challenging due to the need for automation, scalability, and semantic precision across diverse and noisy educational materials. In response to these challenges, we develop a robust and scalable data construction pipeline that transforms raw high school exam papers into structured multimodal instances suitable for training LVLMs, comprising automated problem identification, automated deduplication and diagram extraction, structured data extraction, and manual verification. While the pipeline is largely automated, this lightweight manual verification step is essential for maintaining data quality, particularly when handling complex symbolic expressions and diagrammatic content in real-world settings. Based on this pipeline, we construct AuxSolidMath, a curated dataset of 3,018 solid geometry problems in a rich multimodal format, comprising the problem description, the final answer, the auxiliary-line description, the original diagram, and the auxiliary-line diagram. To our knowledge, AuxSolidMath is the first systematically constructed dataset explicitly tailored to auxiliary-line-based solid geometry reasoning.
@@
We empirically evaluate GeoVLMath against a broad range of LVLMs, including representative approaches that construct auxiliary lines via code-driven rendering. Despite its relatively modest parameter scale, GeoVLMath achieves highly competitive performance, consistently outperforming code-driven rendering methods and surpassing larger models such as Qwen2.5-VL-32B-Instruct (Bai et al., 2025) and GPT-4o (Hurst et al., 2024). These results indicate that supervision grounded in auxiliary-line constructions is more effective for improving geometric reasoning than explicit code-driven diagram rendering or simply scaling model size. Furthermore, our evaluation protocol highlights the value of auxiliary-line-augmented datasets for revealing the limitations of LVLMs in visual reasoning.
Contributions. (1) Cross-modal reward. We introduce a geometry-aware scalar reward that directly evaluates diagram-text alignment, providing reliable and fine-grained supervision for reinforcement learning in auxiliary-line construction. (2) AuxSolidMath. We introduce AuxSolidMath, a curated dataset of 3,018 solid geometry problems from real high-school exams, with aligned diagrams, auxiliary-line annotations, and answers, designed to support training and evaluation of auxiliary-line reasoning. (3) GeoVLMath. We introduce GeoVLMath, demonstrating that RL with vision-based rewards effectively optimizes auxiliary-line construction, achieving superior performance to code-driven rendering approaches and competitive results against larger LVLMs.
@@ 2.1 Problem Definition
We study solid geometry problems whose solutions require the active construction of auxiliary lines. Each instance is defined by a pair $\langle I,q\rangle$, where $I$ denotes the original diagram and $q$ is a question grounded in $I$. The goal is to generate a solution $y$ that includes the auxiliary lines auxaux, a sequence of reasoning steps, and the final answer ansans. The auxaux is not present in $I$ and must be introduced during reasoning to expose implicit spatial relations and convert them into explicit geometric constraints.
Figure 2: Overview of the cross-modal reward-driven RL. We first fine-tune a cross-modal reward model to evaluate the correctness of auxiliary-line constructions. During the RL phase, the reward model’s consistency score is combined with the answer accuracy reward to produce a composite signal that updates the policy via GRPO.
@@ 2.2 Framework Overview
We introduce a two-stage training framework for LVLMs that integrates the auxiliary-line construction into the reasoning process. In the first stage, we apply SFT on automatically synthesized chain-of-thought (CoT) data with explicit auxiliary-line steps, enabling the model to actively construct auxiliary lines, thereby establishing a good initialization. In the second stage, we further use RL to encourage the model to construct auxiliary lines that faithfully reflect the geometry of the diagram, boosting the precision of the solution. At the core is a cross-modal reward model that provides fine-grained feedback by scoring the agreement between the original diagram plus the generated auxiliary-line description and a reference diagram annotated with the correct auxiliary line. In summary, our framework combines direct supervision with structured visual feedback, thereby enhancing the reliability of auxiliary-line constructions and overall geometric problem-solving performance.
2.3 Supervised Fine-Tuning
We apply SFT on CoT exemplars to initialize the model for RL. Auxiliary-line construction steps are explicitly marked with [AUX] and [/AUX] to provide structured supervision and support later reward modeling, and the model is trained using a standard next-token prediction objective.
2.4 Cross-Modal Reward-Driven RL
@@
As shown in Figure 1, accurate auxiliary-line constructions improve reasoning success. The key challenge is how to integrate such constructions into the reasoning process. Recent methods adopt Python-based drawing code to render auxiliary lines for intermediate reasoning, yet their reliance on precise and executable specifications limits robustness in complex solid geometry. Motivated by this limitation, and inspired by recent advances in textual CoT supervision (Xu et al., 2024; Zhang et al., 2025), we represent auxiliary-line constructions as structured natural language descriptions. To bridge the gap between textual descriptions and spatial structure, we propose a cross-modal reward model that scores diagram-text alignment between the original diagram paired with a generated auxiliary-line description and a reference diagram annotated with the ground-truth auxiliary lines. This reward provides geometry-aware supervision without relying on executable drawing code or precise geometric specifications, enabling robust scaling across diverse diagram styles. We integrate this cross-modal reward, alongside a final-answer reward, into a GRPO-based RL stage to align intermediate constructions with the diagram while maintaining final-answer accuracy. An overview of the stage is illustrated in Figure 2.
2.4.1 Cross-modal Reward Model
@@
Given an original diagram $I$, a textual description of auxiliary lines $d$ (either the ground-truth description $d_{\text{aux}}$ or a perturbed variant $d^{-}_{\text{aux}}$), and a reference diagram $I^{+}$ annotated with the ground-truth auxiliary lines, the reward model evaluates the geometric relations implied by applying $d$ to $I$ against the additional geometric structures present in $I^{+}$ but absent from $I$. Rather than relying on surface-level similarity, it assesses diagram-text relational consistency by verifying whether geometric relations specified in $d$, such as parallelism, orthogonality, and angle bisection, are satisfied in the reference diagram $I^{+}$. Accordingly, the proposed reward exhibits three desirable properties: (1) Cross-modal relational alignment, which evaluates diagram-text consistency at the level of geometric relations rather than surface-level or lexical similarity; (2) Render-free and coordinate-free, as it avoids dependence on executable drawing code or pixel-accurate coordinates, relying instead on relational correspondence between diagrams and textual descriptions; and (3) Fine-grained supervision, assigning intermediate scores to partially correct yet geometrically meaningful constructions, thereby enabling precise credit assignment in multi-step reasoning. In summary, the cross-modal reward evaluates diagram-text spatial consistency at the level of geometric relations, assessing whether and to what extent the generated auxiliary lines align with the intended geometric structure, without relying on executable rendering code methods. Next, we describe how to automatically construct diagram-text supervision at scale and use it to train the cross-modal reward model.
@@
Constructing Diagram-Text Supervision. Each training example is represented as $\langle I,d,I^{+},\hat{y}\rangle$, where $\hat{y}=(r,s)$ contains a brief justification $r$ and a consistency score $s\in[0,1]$ indicating how well $d$ aligns with $I^{+}$ given $I$. We construct this supervision dataset via a fully automated pipeline (see Figure 2). Starting from high-quality supervision triplets $\langle I,d_{\text{aux}},I^{+}\rangle$ constructed as described in Section 3, we systematically generate challenging negative descriptions to support robust diagram-text supervision. Specifically, we design a set of rule-based perturbation templates that simulate common auxiliary-line construction errors, including partial deletion, intersection alteration, incorrect line connections, adding irrelevant lines, unrelated auxiliary lines. Building on these templates, we further leverage a large language model to synthesize diverse and linguistically varied negative descriptions $d^{-}_{\text{aux}}$. These negatives are lexically fluent and semantically plausible, yet geometrically inconsistent with the intended construction encoded in the reference diagram $I^{+}$, thereby providing hard counterexamples beyond surface-level mismatches. To assess the consistency between each description $d\in\{d_{\text{aux}},d^{-}_{\text{aux}}\}$ and the target construction $I^{+}$, we adopt an LVLM-as-a-Judge strategy. Concretely, the LVLM is prompted to evaluate diagram-text alignment by comparing the original diagram paired with the description $\langle I,d\rangle$ against the reference diagram $I^{+}$, and to produce both a natural-language rationale $r$ and a scalar alignment score $s\in[0,1]$. This automated evaluation yields interpretable explanations alongside valued supervision signals, enabling scalable and fine-grained diagram-text supervision that spans faithful auxiliary-line descriptions as well as adversarially perturbed constructions.
@@
Training Cross-Modal Reward Model. Given the input triplet $\langle I,d,I^{+}\rangle$, the model outputs a structured prediction $\hat{y}=(r,s)$. We train the model by maximizing the conditional likelihood of the serialized output:
@@
Through this training, we obtain a reward model that provides precise feedback on diagram-text alignment between auxiliary-line descriptions and the reference diagram. This model serves as a key component of our framework, guiding the policy toward auxiliary-line constructions that are geometrically consistent and diagram-grounded.
Figure 3: Overview of the proposed data creation pipeline.